Tidal HI in Merging Galaxies

J. E. Hibbard, Institute for Astronomy

• What is it?

Outer parts of gas-rich disks left behind as the progenitors merge.

• Why is it interesting?

Provides information on the origin/transformation of galaxies — Bears the imprint of the encounter geometry

— Kinematics constrain future evolution of the remnant — Gas content reflects the Hubble type of the progenitors

May have important ramifications — Absorption cross-sections

— Detectability of evolved or distant mergers

— Dwarf formation within tidal tails

• Why am I talking about it?

The gaseous tidal properties of these systems have never been quantified as a class. In the past year we have doubled the number of major mergers mapped in H i (to N=14), and we are now in a position to do just this.

• What will be quantified?

— H i Masses

— Total tidal extent (H i & optical)

— Observability

Observed Systems

As part of two on-going projects investigating the dynamics of major mergers, we have mapped the H i distribution and kinematics in 12 violently interacting/merging galaxies. Data on two more are available from the literature, and similar data will be obtained on a third system shortly. (Two more systems might be available soon from ATNF researchers).

Samples:

• HI Observations of the “Toomre Sequence”

Optically selected sample of relatively isolated galaxies in the NGC catalog with two well defined tidal tail (Toomre,

1977).

— 5 observed by us (Hibbard & van Gorkom 1996)

— 2 from others (Mahoney et al. in prep; English 1995)

• HI Observations of Luminous Infrared Galaxies

We observed the nearest and most luminous galaxies from the IRAS Bright Galaxy Catalog (Soifer et al. 1987).

— All 6 within 100 Mpc with 12 >Log(L_{IR}) >11.5 (5 observed, 1 to come)

— All 3 within 150 Mpc with Log(L_{IR}) >12 (2 detected, 1 not detected)

(Hibbard & Yun in preparation)

See TSeq.gif, ULIR.gif 3

Amounts of Tidal Material

• In a study of 25 optically disturbed systems, Schombert et al. 1990 showed that 15–25% of the total luminousity can reside in the tidal regions. We find similar numbers.

L_{R} ∼ 7 × 10^{9}h_{75}^{−2}L_{⊙}

• We find that from 30–100% of the observed H i can be relegated to the tails. Tidal gas masses are comparable to the total gas mass of a normal spiral.

M_{HI} ∼ 4 × 10^{9}h_{75}^{−2}M_{⊙}

• H i only associated with low velocity dispersion components (tails and disks). Plumes and envelopes are H i-poor.

(drew intensity profiles for each)

• In several cases, extensive H i tails were discovered where there were no known optical counterparts. Some low surface brightness optical counterparts have subsequently been imaged, but some seem devoid of starlight to μ_{R}>27 mag arcsec^{−2}.

See HImass.ps. Also showed mass histogram from Haynes & Giovanelli 1984, AJ, 89, 758

Extent of Tidal Features

• Often reported that disturbed galaxies subtend a much larger cross-sectional area than normal galaxies.

— M51 system (Rots et al. 1990, AJ, 100, 387)

length ∼90 kpc

— M81 system (van der Hulst 1979, A&A, 75, 97) diameter ∼100 kpc

— NGC 2444/5 (Appleton et al. 1987, Nature, 330, 140) length ∼200 h_{75}^{−1} kpc

— Mkn 348 (Simkin et al. 1987, Science, 235, 1367) diameter ∼280 h_{75}^{−1} kpc

• Since D_{HI} > D_{opt} in normal spirals, is D_{HI} >> D_{opt} in mergers?

See CrossSect.gif & ImpactParam.gif 7

Extent of Tidal Features (conclusions):

• While mergers appear to lie in the upper range of the area covering factor plots (perhaps due in part to the fact that two systems have contributed to the tidal debris), the increase is not an order of magnitude larger than field spirals.

— material becomes fainter as it is pulled out

• However, the maximum radial extent is considerably larger

— resulting covering factors are ∼0.2–0.4 in outer regions

• The optical light is not that much less widely distributed than HI at the faintest measured levels.

— H i might actually be less widely distributed with respect to the optical light as compared to field galaxies.

Ionization of outer HI?

Higher velocity dispersion of stars?

Further Extent of Tidal Features

HI Tails just represent the highest phase-space density (highest density, lowest velocity dispersion) tidal material. This will be the proverbial “tip of the iceberg” of a much more pervasive lower phase-space density medium.

• Lower N_{HI} former disk material, possibly ionized

• Low density, high dispersion halo material

Observed tidal extent is a lower limit to the true extent of these systems

See Mice.gif 10

Future Evolution and Continued Observability

• Detailed simulation of the NGC 7252 merger remnant (Hibbard & Mihos 1995) suggests that the majority of tidal material remains bound to the remnant, with the outer regions on very long-period orbits.

— Short, short-lived, luminous tails yield to extensive, longlived, low surface brightness tails.

— This material evolves on very long timescales, and should remain observable for several more Gyrs.

— In the absence of external forces (intruding galaxies, cluster potential, IGM), this material will extend to very low column densities at very large distances from the remnant.

• Returning tidal material creates higher phase-space density features as it wraps coherently within the potential (e.g., Hernquist & Spergel 1992). Should leave visible features for 2–5 Gyr.

See EarlyEvol.gif & LateEvol.gif 12

Observability in Redshift Space

Simulate observation of 5 mergers at z=0.3, z=0.5 (q_{o}=0, H_{o}=100) with CFHT 3.6m telescope

• Surface brightness fades as (1 + z)^{−4}

z = 0.01 → 0.1 0.3 0.5 μ = 25.0 B → 25.4 B 26.1 V 26.8 R

• Angular size decreases as (1 + z)^{2}/(2z + z^{2})

z = 0.01 → 0.1 0.3 0.5 θ = 2′ → 30′′ 13′′ 9.5′′

(Euclidean space: →  26′′  9′′ 5.3′′)

• Scale sky background according to observing band

Moderate brightness tidal features are recoverable from the ground with a 1hr exposure at z=0.3, and a 4hr exposure at z=0.5 with a 4m class telescope.

See lowZ.gif, highZ.gif 14

General Kinematic Properties of Merging Galaxies:

The majority (10/14) conform (at least to first order) to the simple tidal dynamics expected for merging galaxies: tidally expanding streams of gas and stars with smooth runs of velocity, connecting smoothly back to the progenitors or remnants. (Arp 295, IC 883, Mkn 273, VV114, NGC 2623, NGC 3256, NGC 4038/9, NGC 4676, NGC 3921, NGC 7252)

The remaining four are quite baffling:

• NGC 520: Rotating H i disk cut by stellar streams.

• Arp 299: Long H i plume offset from stellar tail.

• Arp 220: Outer H i anti-correlated with stellar features (similar to NGC 520?)

• NGC 1614: Outer H i tail does not connect obviously back to stellar tail

Conclusion:

Non-gravitational forces are playing an important role in the dynamics of some merging systems.

See N520.gif, A220.gif, A299.gif 16

Tidal Dwarfs

The propensity for dwarf-sized clumps to appear within tidal filements was noted by Zwicky (1956), and presented in some detail by Schweizer (1978).

Roughly Three Flavors:

• Detached Stellar Systems Entrained within a connecting HI stream

MW & LMC-SMC: Mathewson et al. 1974 M81 & NGC 3077: van der Hulst 1978; Yun 1992 NGC 520 & UGC 957: Stanford & Balcells 1991; Hibbard & van Gorkom 1993 Arp 105: Duc & Mirabel 1994

• dIrr–sized blobs within stellar & HI tail, often actively forming stars

The Antennae (NGC 4038/9): van der Hulst 1978; Schweizer 1978; Mirabel, Dottori & Lutz 1992 NGC 7252: Schweizer 1978; Hibbard & van Gorkom 1993 NGC 3921: Schweizer 1978; Hibbard et al. 1994

• Optical “Knots” in Tidal Features

The Superantennae: Mirabel, Lutz, & Maza 1991 in IRAS galaxies: Hutchings 1996 in HCGs: Hunsberger, Charlton & Zaritsky 1996

While integrated properties are consistent with those of legitimate dwarfs, better kinematic evidence that systems are indeed selfgravitating is sorely needed.

See N7252dw.gif, N3921dw.gif, N4676.gif 18

Tidal HI in Merging Galaxies CONCLUSIONS:

• Most mergers have tail kinematics consistent with simple tidal models (Toomre & Toomre 1972). Tails contain ∼ 7 × 10^{9}h_{75}^{−2}L_{⊙} of starlight and ∼ 4 × 10^{9}h_{75}^{−2}M_{⊙} of neutral gas.

• Luminous tails raised initially, but most material falls back quickly. Low surface brightness features should persist for several Gyr. Actual timescales will depend on encounter geometry.

• Mergers present a large cross-sectional area, although not orders of magnitude larger than normal spirals at a given isophote. However, they DO have high surface brightness/column density material at much larger radii than normal spirals.

• There must be more extensive fainter material beyond what is observed

• “Classical” mergers are observable to z ≈ 0.3–0.5 with deep ground-based imaging

• Star-forming clumps of gas and stars are often found within the expanding tidal tails. These concentrations have global properties similar to dwarf irregular galaxies, and may evolve into long-lived companions to the remnants.

• We don’t understand everything yet — some very intriguing differences between gaseous and stellar tidal features in some systems